OpenStudy (anonymous):

Solve: 4(x − 1) − (x + 2) = 12

OpenStudy (precal):

distribute 4 and then your -1

OpenStudy (anonymous):

@precal thanx

OpenStudy (anonymous):

4x-4-x-2=12 3x-6=12 3x=18 x=18/3 x=6

OpenStudy (anonymous):

@Yashhere thanx

OpenStudy (anonymous):

@terrell352 my medal please

OpenStudy (precal):

@terrell352 that is another way to thank someone, just award the medal for who you think gave you the best answer.